A capacitor of capacitance C is charged to a potential difference V from a cell and then disconnected from it . A charge +q is now given to its positive plate. the potential difference across the capacitor is

Grade:12th pass

2 Answers

Khimraj
5 years ago
initial condition
Q = CV
V = Q/C
final condition 
charge will redistribute and plates have charge Q + q/2 and -Q – q/2
so now V = (Q + q/2)/C = Q/C + q/2C = V + q/2C
Hope it clears.
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your question.
 
Just after the cell is disconnected, the capicitor will have
Q = CV
V = Q/C
When +q charge is given to positively charged plate, it will redistribute and the plates would acquire the charges Q + q/2 and -Q – q/2 respectively.
Hence, The new potential difference across capacitor, V’ = (Q + q/2)/C = Q/C + q/2C = V + q/2C
 
Hope it helps.
Thanks and regards,
Kushagra